### v1

# Carrer Pic de Peguera 13, 17003 Girona, Spain

Abstract
The combined use of event-based vision and Spiking Neural Networks (SNNs) is expected to significantly impact robotics, particularly in tasks like visual odometry and obstacle avoidance. While existing real-world event-based datasets for optical flow prediction, typically captured with Unmanned Aerial Vehicles (UAVs), offer valuable insights, they are limited in diversity, scalability, and are challenging to collect. Moreover, there is a notable lack of labelled datasets for underwater applications, which hinders the integration of event-based vision with Autonomous Underwater Vehicles (AUVs). To address this, synthetic datasets could provide a scalable solution while bridging the gap between simulation and reality. In this work, we introduce eStonefish-scenes, a synthetic event-based optical flow dataset based on the Stonefish simulator. Along with the dataset, we present a data generation pipeline that enables the creation of customizable underwater environments. This pipeline allows for simulating dynamic scenarios, such as biologically inspired schools of fish exhibiting realistic motion patterns, including obstacle avoidance and reactive navigation around corals. Additionally, we introduce a scene generator that can build realistic reef seabeds by randomly distributing coral across the terrain. To streamline data accessibility, we present eWiz, a comprehensive library designed for processing event-based data, offering tools for data loading, augmentation, visualization, encoding, and training data generation, along with loss functions and performance metrics.
keywords:
spiking neural networks, event-based sensor, synthetic datasets, optical flow, underwater datasets, event-based cameras, events augmentations
1 Introduction
Event-based Cameras (EBCs), also known as dynamic vision sensors or neuromorphic cameras, are imaging sensors that operate differently from traditional frame-based cameras. Inspired by the exceptional motion perception abilities of winged insects, these cameras respond asynchronously to brightness or intensity changes with microsecond resolution (Lichtsteiner et al., 2008; Brandli et al., 2014). EBCs offer significant advantages over traditional frame-based cameras: no motion blur, high dynamic range, high temporal resolution, and low latency (Gallego et al., 2020). These unique properties make the EBC an ideal sensor to analyse dynamic scenes characterized by fast motion and rapid changes in lighting conditions.
EBCs and optical flow prediction are two complementary concepts that play an integral part in motion analysis in computer vision. Traditionally, optical flow prediction has relied on analysing a sequence of frames to infer motion patterns (Shah and Xuezhi, 2021). However, EBCs present a new paradigm by providing an asynchronous and sparse representation of data in the form of a stream of events (Gallego et al., 2020). As such, many methodologies have been developed for event-based processing, though primarily for relatively simple scenes containing well-defined shapes, edges, and corners (Kueng et al., 2016; Zhu et al., 2017; Gehrig et al., 2020). However, these methods have proven to be inefficient and computationally expensive in complex scenes.
An alternative approach is the use of Neural Networks (NNs) for processing event-based data, which has demonstrated remarkable success in modelling and understanding the nature of these biological vision systems. One significant drawback of EBCs, however, is their incompatibility with conventional NN models, such as Convolutional Neural Networks (CNNs). This requires the use of complex data encoding schemes to convert the sparse asynchronous stream of events into a frame-based representation that can be handled by CNNs, as demonstrated by researchers (Zhu et al., 2018b, 2019; Wan et al., 2022) who have implemented U-net (Weng and Zhu, 2015) like architectures for event-based optical flow estimation. However, such compression of event streams into frame-based representations results in loss of information, thereby limiting the potential of EBCs. As a result, there is a demand for newer paradigms specifically optimized for handling such data. An SNN is one such brain-inspired computing model that relies on so-called spiking neurons to process information (Tavanaei et al., 2019). Spiking neurons are inherently designed to process temporal patterns, making them naturally compatible with the asynchronous stream of events. An interconnection of such spiking neurons in an SNN allows them to effectively extract spatial features, enabling a comprehensive framework for encoding dynamic visual motion over time.
Nonetheless, there is still a lack of sufficiently diverse and well-labelled datasets to facilitate model training. Generating real-world data is both expensive and time-consuming, often requiring specialized equipment to obtain corresponding ground truth data as an alternative to the tedious process of manual annotations by human experts. For instance, datasets such as MVSEC (Zhu et al., 2018a) and DSEC (Gehrig et al., 2021), utilize 3D LiDARs to create a 3D map of the environment. This reconstructed environment is then projected onto the EBC sensor to generate a dense optical flow for use as ground truth during training. However, the resulting optical flow includes some inaccuracies due to clock synchronization issues and often tends to miss information, failing to fully represent the event sequences captured by the sensor. As such, these real-world datasets may not be ideal for model training or evaluation, which stresses the need for a synthetic alternative. In fact, due to the nature of EBCs, synthetic event-based data may translate well to real-world data. The features extracted from synthetic frame-based camera datasets often exhibit obvious differences from the features extracted from real-world data. These differences typically arise from distinct lighting conditions, camera noise, and environmental complexity in real-world settings which are challenging to replicate accurately in synthetic data. Such variations can lead to shifts in pixel intensities, colour distribution and texture details, resulting in significant differences between the features extracted from synthetic and real-world images. In contrast, EBCs only trigger events when detecting brightness changes, producing an image of events that only has 2 values: a positive or a negative polarity. As such, unlike frame-based images that rely on a static grid of pixels with values typically ranging between 0 and 255, the output of EBCs is simpler, leaving more room for error when transitioning from simulation.
To our knowledge, no event-based optical flow prediction datasets currently exist for underwater environments. Capturing such data in real-world conditions presents significant challenges due to several factors such as poor lighting conditions, water turbidity, and light scattering. These factors adversely affect the performance of conventional camera systems typically used for capturing optical flow data. To overcome these limitations, we take a synthetic approach, enabling precise control over environmental conditions and ensuring reliable ground truth. A synthetic event-based dataset for underwater optical flow prediction opens new avenues for research on EBCs in these conditions. Light absorption and scattering, especially at greater depths, reduce illumination and cause colour distortion. Unlike conventional cameras, EBCs operate based on changes rather than ambient light, allowing them to function effectively in low-light scenarios. Moreover, EBCs are highly resilient to motion blur and can effectively filter out noise caused by scattering effects, making them ideal for capturing fast-moving objects in dynamic underwater environments. Their high dynamic range further enables them to overcome the contrast limitations of conventional frame-based cameras.
As such, we present eStonefish-scenes, a synthetic event-based optical flow dataset tailored specifically for underwater environments, together with an easy-to-use pipeline for creating custom datasets. eStonefish-scenes is based on the Stonefish simulator (Grimaldi et al., 2025), with a Remotely Operated Vehicle (ROV) capturing event-based, grayscale, and optical flow data. To further enhance the realism of our simulations, we introduce Stonefish-scenegen, a scene generator capable of randomly populating the environment with corals, creating rich and vibrant life-like reefs. To animate these scenes, we also present Stonefish-boids, a package which simulates the behaviour of schools of fish, adding realistic dynamics to the environment. Together, these components enable the creation of highly realistic underwater scenarios. The dataset is hosted on Zenodo (10.5281/zenodo.15130453) and the data generation pipeline is publicly available at the following link: https://github.com/CIRS-Girona/estonefish-scenes.
Finally, we also introduce eWiz, a library that is composed of all the tools necessary for processing event-based data, ranging from data loading and manipulation, to augmentation and visualization, to encoding and training data generation, together with various loss functions and performance metrics. We plan to add more functionalities and incorporate more areas of applications as our work on event-based processing progresses. The library is openly available at the following link: https://github.com/CIRS-Girona/ewiz.
In summary, our main contributions are:
  • The creation of eStonefish-scenes, an event-based optical flow dataset specifically tailored for underwater environments, along with an open data generation pipeline, which simplifies the process of extending the datasets. To the best of our knowledge, no prior event-based datasets exist for underwater applications.
  • The development of Stonefish-scenegen, a scene generator for Stonefish to streamline creation of diverse, coral-rich underwater environments.
  • The development of Stonefish-boids, a package for simulating schools of fish in Stonefish, enabling the generation of dynamic underwater scenes.
  • The development of eWiz, a comprehensive framework for event-based data processing and manipulation.
The remainder of this paper is organized as follows. Section 2 presents an overview of other publicly available datasets for event-based optical flow prediction. Section 3 and Section 4 describe the details of the simulation setup and sensor configurations used for the dataset, along with the data generation pipeline. Section 5, on the other hand, presents details on the eWiz library for additional processing of the datasets for visualization, analysis and model training. Section 6 gives an overview of how the datasets are structured. Finally, Section 7 concludes this study and presents directions for future work.
2 Related Work
In the literature, only a limited number of real-world event-based datasets have been employed for developing and evaluating event-vision algorithms and NNs targeting optical flow prediction. Among these, the widely used MVSEC dataset (Zhu et al., 2018a) stands out, offering two distinct scenarios: a stereo pair of EBCs mounted on an UAV in an indoor environment, and the same setup mounted on a car in an outdoor setting. The ground truth optical flow for indoor scenes is derived using a motion capture system, while for outdoor scenes, it is obtained through a combination of GPS data and scene reconstruction techniques. Building upon MVSEC, the DSEC dataset, introduced by Gehrig et al. (2021), presents significant improvements. It features driving scenarios with displacement field magnitudes reaching up to 210 pixels and a camera resolution three times higher than that of MVSEC. While these datasets have been instrumental in advancing event-based vision research, they are not without limitations. The generation of real-world data is both resource-intensive and time-consuming, often necessitating specialized equipment to capture accurate ground truth information. This reliance on costly setups further underscores the challenges of producing high-quality datasets in this domain. Moreover, the MVSEC dataset exhibits notable limitations in the accuracy of its ground truth optical flow. The ground truth was generated using a Velodyne LiDAR, which relies on a 3D reconstruction of the environment to estimate optical flow. The dataset suffers from additional challenges due to the mechanical setup. Specifically, the vibrations of the car during forward motion induce a wobbling effect in the event-based sensor. This wobble results in the sensor capturing significant up-and-down motion, whereas the ground truth optical flow predominantly reflects forward motion. This mismatch between the motion perceived by the event-based sensor and the ground truth optical flow can introduce discrepancies in training and evaluating models, especially for applications that rely on precise motion estimation.
These challenges contribute to the scarcity of real-world event-based datasets, prompting researchers to increasingly turn to synthetically generated datasets. Leveraging advanced simulators, synthetic datasets provide a scalable and cost-effective alternative, enabling the generation of diverse and accurately labelled data. ESIM (Rebecq et al., 2018) for example, being one of the most accurate in the literature, makes use of an adaptive rendering scheme that only samples frames when necessary. Such a rendering scheme adapts the sampling rate to the predicted dynamics of the visual signal, allowing for a faithful simulation of events when fast motion occurs. ESIM has been used to generate several synthetic event-based optical flow datasets, including MultiFlow (Gehrig et al., 2024), and HREM (Luo et al., 2024). MultiFlow contains multiple moving objects undergoing continuous similarity transformations, along with dense ground-truth optical flow. HREM, on the other hand, comprises 100 virtual scenes designed to replicate real-world indoor and outdoor environments. BlinkSim, proposed by Li et al. (2023), is another EBC simulator that includes a powerful rendering engine capable of building thousands of scenes with different objects and motion patterns at high frequency. This simulator gave rise to the BlinkFlow dataset, which improves generalization performance of state-of-the-art methods by 40% on average. Another synthetic dataset is the MDR created by Luo et al. (2023). It makes use of Blender to create indoor and outdoor 3D scenes and includes diverse camera motions and high frame-rate videos between images.
One of the reasons behind generating synthetic datasets is the ability to diversify the created environments and easily extend applications related to EBCs. Although synthetic optical flow datasets for event-based vision are already available, most are not specifically designed to reflect the motion dynamics or characteristics of any particular vehicle type. MDR, ESIM, and BlinkFlow are limited to predefined camera trajectories, which restricts their ability to accurately replicate such dynamic motion patterns. Furthermore, these simulators lack sophisticated light behaviour modelling, which is essential for creating realistic and complex environments. For example, factors such as absorption, scattering, and refractive effects, key characteristics of underwater environments, that these simulators fail to account for, undermine the fidelity of synthetic datasets for underwater applications. On the other hand, ESIM makes use of an OpenGL rendering engine (The Khronos Group, Inc., 2014). Similarly, the use of Blender The Blender Foundation (2002) for MDR requires extensive expertise in 3D scene creation to appropriately simulate real-world lighting conditions, especially for underwater environments. All these factors, make the creation of synthetic event-based datasets a challenging endeavour, particularly for underwater environments.
Currently, research on underwater event-based applications is highly limited, especially due to the lack of publicly available datasets captured with EBCs in underwater environments. To address this gap, we created a diverse synthetic dataset tailored for underwater applications, along with a modular data generation pipeline to facilitate creation of custom datasets. The dataset incorporates both static and dynamic elements, enabling the trained algorithms to learn robust feature representations by capturing background stability and motion-driven variations. This would also enhance generalization to real-world scenarios, where stationary structures and moving entities, such as marine life, coexist.
Finally, although a few open-source tools are available for specific event-based data processing tasks, such as the work by Shiba et al. (2022), no comprehensive library currently integrates these capabilities into a unified framework. To address this gap and streamline future development, we developed eWiz, a comprehensive library for event-based data manipulation and processing.
3 Simulation Setup
This section describes the simulation environment, sensor and vehicle configurations, and the data collection approach for generating the event-based optical flow prediction dataset, eStonefish-scenes, tailored for autonomous underwater vehicles. The dataset consists of event streams, grayscale images, and the corresponding ground truth optical flow. Additionally, it utilizes an optimized format provided within the eWiz framework to facilitate efficient data storage, access, and processing. Further details on this optimized format are provided in Section 5.
The dataset was generated using the Stonefish simulator (Grimaldi et al., 2025) within a custom-designed environment featuring a texture-rich seabed. To add visual complexity, the seabed was populated with abundant corals and marine flora with pronounced edges and contrasts. The corals were divided into 3 clusters distributed across the seabed, with each cluster containing 20 corals of different sizes, ranging from small antler corals to large dome corals. The scene was randomly generated using a scene generator to diversify coral pose and size. The dataset features two environments:
  • Rocky: Datasets with the rocky environment capture a rocky seabed texture. It is characterized by a barren rocky seabed with no flora in the scene.
  • Reef: Datasets with the reef environment are characterized by the same rocky seabed texture, but populated with flora to simulate the look and feel of a colourful coral reef.
The dataset comprises both, static and dynamic scenes. Static scenes are simple, empty environments with no dynamic objects, such as schools of fish. Only stationary corals appear in these scenes, ensuring consistency with event-based optical flow loss functions commonly utilized in the literature (Stoffregen and Kleeman, 2019). Dynamic scenes, on the other hand, incorporate dynamic elements in the form of schools of fish. Several fish models were used across the simulation, each exhibiting slightly modified Boid behaviours (Reynolds, 1987) to mimic the behaviour of schools of fish in nature. This diversity in scene complexity, creating complex visual patterns, provides a comprehensive framework for both training and evaluating optical flow algorithms across various environments. Figure 1 depicts an example of the generated scene.
Figure 1: Generated environment, the environment includes a seabed with feature-rich corals
For data collection, a joystick-controlled BlueROV2 (Blue Robotics, 2016) model was utilized as shown in Figure 2. The ROV was equipped with an RGB camera, an event-based camera, and an optical flow sensor, mounted in two configurations. To ensure diverse and robust data, the scene was recorded from multiple viewpoints, mimicking the placement of a camera on a real ROV. Specifically, the sensors were positioned in the following configurations:
  • Forward-looking: In this setup, the depth of the ROV was adjusted dynamically while moving.
  • Down-looking: In this setup, the ROV maintained a constant depth while moving, with a focus on the seabed.
The forward-looking configuration is particularly advantageous for tasks such as scene inspection and obstacle avoidance, as it provides a clear view of the environment ahead. In contrast, the down-looking configuration is optimized for computing the odometry of the robot, leveraging the EBC feed to analyse the textured seabed. These complementary configurations enable the ROV to address a diverse range of underwater exploration and navigation challenges.
Figure 2: Simulated BlueROV2 navigating an underwater environment.
[…]
Sequence Number Scene Type Altitude Environment Camera Position
1 Static Constant Reef Down-looking
2 Static Varying Reef Front-looking
3 Static Constant Rocky Down-looking
4 Static Varying Rocky Front-looking
5 Dynamic Constant Reef Down-looking
6 Dynamic Varying Reef Front-looking
7 Dynamic Constant Rocky Down-looking
8 Dynamic Varying Rocky Front-looking
Table 1: Summary of the generated sequences and their corresponding configurations.
These sequences can be utilized for training neural networks by dividing them into separate event volumes. Each event volume contains a stream of events, along with all grayscale images that occurred at the corresponding timestamps. One such example is by extracting events that took place between every pair of consecutive grayscale images. The resulting volumes would include a grayscale frame at time $t$, another frame at time $t+1$, and all events that happened in between. Based on the 20 FPS refresh rate of the grayscale sensor, we can extract 800 training samples per sequence approximately, providing around 6400 samples for all sequences combined.
4 Data Generation Pipeline
The dataset can be easily extended by using a modular data generation pipeline, which comprises 3 separate packages, each responsible for different aspects of data generation. First, we present the Stonefish-scenegen package. This package is used to generate coral-rich and vibrant scenes by randomly populating the seabed with corals of different types and sizes. It generates a Stonefish-compatible scene file, which can then be used to launch the main pipeline, eStonefish-scenes. The pipeline is a ROS package with a launch file for Stonefish. The launch file starts the simulator with the generated environment, spawning the ROV and displaying live feeds from all its sensors. The pipeline internally uses the Stonefish-boids package to randomly spawn schools of fish in the scene. The user can then move the ROV around the scene while recording a rosbag, which can subsequently be converted to the eWiz format using the ConvertStonefish module of eWiz.
[…]
4.1 Scene Generation
One challenge with the Stonefish simulator is the time-consuming process of creating rich and dynamic underwater scenes. Each coral or 3D object must be manually defined in a scene file, specifying its pose in the world frame, size, and physical attributes. While this approach is feasible for a few objects, it becomes impractical for environments featuring numerous objects. To address this limitation, we introduce Stonefish-scenegen, an underwater scene generator specifically designed for the Stonefish simulator. This tool streamlines the creation of complex underwater environments, significantly reducing the effort required for scene generation. It consists of a Python script that takes as input a Stonefish scene file, a ground surface, and the desired 3D models to spawn. The user chooses the number of coral clusters, along with the number of corals spawned in each cluster. Each coral instance is placed by sampling a random vertex on the ground surface. The normal of the surface passing through that vertex is then used to determine translation and rotation transformations. To ensure reproducibility, the user can generate the exact same scene by fixing a seed before running the script. The script also converts from the ENU coordinate system to its NED equivalent, as required by Stonefish. Finally, a scene file containing the path, texture, size, and pose of all spawned 3D models is generated. This scene file can then be used with eStonefish-scenes to launch the simulation. The scene generator is made openly available at the following link: https://github.com/CIRS-Girona/stonefish-scenegen.
4.2 Simulating Schools of Fish
To enhance the visual complexity of the underwater scenes with some dynamic elements, we incorporated an adapted implementation of the Boids algorithm (Reynolds, 1987) for simulating the natural grouping behaviour observed in schools of fish. The algorithm is based on four fundamental rules that govern the behaviour of each fish within the school, collectively influencing the behaviour of the group as a whole.
  • Alignment: Each fish adjusts its velocity to match the average velocity vector of nearby fish within a defined perception radius.
  • Cohesion: Each fish steers toward the average position of nearby fish within its perception radius.
  • Separation: Each fish receives a repulsion force if it gets too close to another fish in the school. This helps maintain a minimum distance between all fish.
  • Leader Attraction: One fish is designated as the leader of the school and follows a predefined path along the map, while the other fish experience an attractive force between them and the leader.
At each simulation step, the total force acting on each fish is calculated as a weighted sum of the contributions from each rule, which is then used to derive the change in position. This force-based approach can be extended beyond the four fundamental behavioural rules, allowing additional forces to dynamically influence movement in response to the surroundings. For instance, external forces such as repulsion from obstacles help maintain realistic movement patterns by preventing collisions with elements like corals, miscellaneous structures, the water surface and the ROV.
To efficiently handle such environmental interactions with static objects, we use OctoMap (Hornung et al., 2013) for discretizing the simulation space, as illustrated in figure 4. This enables rapid detection of potential collision zones and the assignment of repulsive force vectors to different locations. The flexibility of the library allows for the creation of a custom Octree where each node stores a repulsive force vector normal to the object surface. When a fish enters an obstacle voxel, a repulsive force is applied, ensuring smooth and natural avoidance behaviour. For dynamic objects such as the ROV, on the other hand, repulsive forces are applied based on the distance between each fish and a defined repulsion radius around the object, allowing for real-time interaction.
(a) Octree Scenario
(b) Seabed and Corals Scenario
Figure 4: Discretized underwater environment as an Octree structure.
Additionally, each school of fish follows a randomly generated path created using the Rapidly-exploring Random Trees (RRT) algorithm (LaValle, 1998), implemented through the Open Motion Planning Library (OMPL) (Şucan et al., 2012). The designated leader follows this path, while the rest of the group moves along with it based on the leader attraction rule. Each time a new simulation is initiated, both the path and the spawning positions of the fish are randomly generated. However, for reproducibility, users can set a fixed seed to ensure the path and spawning positions remain consistent across runs.
Users can customize the type of fish schools and adjust various behavioral parameters through a YAML configuration file. These include alignment, cohesion, and separation weights, as well as spawning location, model scale, and school size. This functionality is provided as an open-source ROS package, Stonefish-boids, which builds on the base Stonefish API. The package is available at: https://github.com/CIRS-Girona/stonefish_boids.
4.3 Data Collection
To record data from the generated scenes, we provide a ROS-based utility, eStonefish-scenes, which uses the scene file generated by the scene generator and internally calls stonefish-boids to simulate schools of fish. It relies on the Stonefish-ROS bridge to facilitate communication with the simulator. The package also includes meshes and textures for the BlueROV2 and a rocky seabed. A predefined scenario combines these assets into a single environment, spawning the ROV and the seabed beneath. A Python script is provided to control the robot, utilizing its allocation matrix to enable joystick-based operation. Additionally, an events rendering script is included to visualize the resulting event image in RViz (Kam et al., 2015) as the simulation progresses. Users can customize the sensors mounted on the ROV by modifying the corresponding scene file. Customization options include, but are not limited to, adjusting resolution, field of view, event-based sensor thresholds, and adding noise. The record.bash script allows users to record rosbags from the simulation, which can subsequently be converted to the eWiz format using the ConvertStonefish module of eWiz. This data collection utility is available at the following link: https://github.com/CIRS-Girona/estonefish-scenes.
5 Data Processing
The eStonefish-scenes dataset was generated with NN training and inference in mind. To streamline this process, we created eWiz, a Python-based library that can be installed with PyPi (The Python Community, 2003), serving as an all-inclusive framework for event-based data manipulation, processing, visualization, and evaluation. eWiz is designed for seamless integration with libraries such as PyTorch (The PyTorch Foundation, 2016) and Tonic (Gregor Lenz, 2019), making it suitable for ML-oriented pipelines. Additionally, it offers compatibility with existing datasets and supports both event-based and frame-based data. The following subsections present an overview of the various functionalities currently offered by eWiz. For further details, the readers are encouraged to refer to the official documentation. The library is openly available at the following link: https://github.com/CIRS-Girona/ewiz.
Figure 5: Currently implemented modules for the eWiz library.
5.1 Data Access and Storage
Due to the high temporal resolution and data throughput of EBCs, event-based data requires substantial storage space. Additionally, slicing data between any two timestamps requires time-consuming search algorithms due to the large volume of events in the sequence. Libraries like NumPy Harris et al., 2020 also face high memory demands, as they typically require loading entire sequences into memory before slicing or processing. The eWiz format addresses these challenges with an optimized approach to data management and manipulation. Firstly, we use Blosc (The Blosc Developers, 2014), a high-performance compressor optimized for binary data, effective for compressing numerical arrays. To further reduce data size, we constrain the possible values for every event stream. For instance, pixel coordinates are stored as 16-bit unsigned integers, while polarities are stored as boolean values. Next, we use the h5py library (Andrew Collette and Contributors, 2017) read and write data as HDF5 files (The HDF Group, 2006), which supports large and complex data structures, and allows slicing without the need to completely load the data into memory. Finally, to reduce loading times, we pre-compute the mappings between timestamps and data indices during the data writing process. This includes binary search to map timestamps (in $ms$) to event, grayscale, and flow indices, as well as direct mappings of event indices to their grayscale and flow counterparts. Additionally, optical flow data is automatically synchronized between specified timestamps by interpolating and accumulating flow displacements as needed.
eWiz provides easy-to-use readers and writers to facilitate this process. The data reader uses the precomputed time mappings to read the desired indices of events based on the start and end timestamps and loads them to memory. It also loads the corresponding grayscale images that were captured at the beginning and at the end of the event stream along with the synchronized optical flow. The user has the possibility to stride over the events, by either using event indices, timestamps, or grayscale images indices. The data reader goes hand in hand with the included data iterators which allow to sequentially stride over the data. Such functionality allows the user to obtain event-based data sequentially, without the need for manual indexing.
eWiz also aims to support a multitude of EBCs and ensure compatibility with publicly available datasets. We include several data converters for transforming these data types into the eWiz format. Currently, we support the DAVIS346 (Inivation, 2019) and Prophesee (Prophesee, 2024) cameras, along with popular datasets such as MVSEC (Zhu et al., 2018a) and DSEC (Gehrig et al., 2021).
5.2 Data Encoding and Augmentation
Traditionally, CNN-based approaches (Zhu et al., 2018b, 2019; Wan et al., 2022) require the EBC output to be converted from raw streams of events data to 2-channel event images, where one channel represents positive polarities and the other the negative polarities. To support this, eWiz provides different encoders. Currently, two encoding schemes are implemented: Gaussian (Ding et al., 2022) and event counting Nguyen et al. (2019). The former computes pixel values by summing each event’s polarity weighted by a normalized Gaussian kernel based on the event timestamp and a scaling factor $\lambda$ for each bin. Whereas, the latter simply groups events into fixed time intervals, with pixel values computed by summing the polarities of all events within each bin.
eWiz also provides numerous event-based temporal and spatial data augmentations, including techniques such as time warping, noise injection, event flipping, and spatial transformations, all of which are fully compatible with PyTorch and Tonic. These augmentations enable enhanced model generalization by simulating diverse real-world conditions, making them ideal for training robust event-based neural networks.
5.3 Data Visualization
eWiz includes several rendering modules for visualizing the acquired data. The primary renderer replays captured sequences, overlaying event images on top of grayscale images for an intuitive view of the data. Additionally, the library offers separate visualizers for event-based data, grayscale images, and optical flow, with the capability to save rendered sequences. Users can visualize the data as video sequences, individual (encoded) frames, or 3D volumes of event streams with interleaved frames.
5.4 Algorithms
eWiz not only supports data processing for NN-based approaches but also algorithms, such as contrast maximization (Shiba et al., 2022), that directly operate on raw event streams. Contrast maximization involves warping the event stream based on a proposed optical flow, iteratively optimizing the optical flow field to maximize the contrast computed between the warped and the original event streams. Higher contrast, indicated by well-defined edges, signals improved alignment of events. As such, eWiz also provides warping tools that allow the implementation of various motion models based on the camera movement type. In the long run, eWiz will support not only optimization-based algorithms but also deep learning-based and SNN-based alternatives. We plan to add more functionalities, such as ground truth generation via optimization and SNN-based training algorithms.
5.5 Loss Functions and Evaluation Metrics
eWiz supports multiple loss functions for training CNNs and SNNs, currently focused on optical flow prediction. For instance, it supports the photometric loss (Zhu et al., 2018b), as a self-supervised approach to minimize pixel-wise intensity differences between two grayscale images encompassing an event stream. The predicted flow field is used to warp the grayscale images and optimize alignment. This loss can be combined with Charbonier loss (Sun et al., 2014), a smooth approximation of the L1 loss that is more sensitive to large outliers, as well as a smoothness loss serving as a regularizer to address the aperture problem and ensure continuous flow across sparse event-based data. In addition, the library includes losses specifically designed for the contrast maximization algorithm (Gallego et al., 2018, 2019) to evaluate warped event streams. These include image variance and gradient magnitude loss functions, along with their improved variants presented by Shiba et al. (2022), where the losses are normalized and applied in a multi-focal manner. To further enhance performance, eWiz also provides a regularizer that smooths the predicted flow across the image, helping address the sparsity of event-based data and ensuring continuous flow, ultimately improving convergence during training.
Moreover, eWiz can be used not only to evaluate the performance of trained models but also to track their performance during training. Currently, the library supports performance metrics for optical flow prediction tasks, such as the Average Endpoint Error (AEE) (Zhu et al., 2018b), which is defined as the distance between the endpoints of the predicted and the ground truth flow vectors, and the Average Angular Error (AAE) (Zhu et al., 2018b), defined as the angular difference between the prediction and ground truth flows, both averaged over the number of pixels. Due to the sparsity of event-based data, these metrics are only computed for pixels where events have occurred. Another implemented metric is the outlier percentage, which refers to the percentage of pixels in the image with an endpoint error higher than a certain value. We define XPE the overall outlier percentage of pixels with endpoint errors greater than X.
6 Dataset Structure
The eStonefish-scenes dataset follows the eWiz format, which stores data in HDF5 files (The HDF Group, 2006). Specifically, events.hdf5 holds the event streams, gray.hdf5 contains grayscale image sequences associated with each event stream, and flow.hdf5 stores the optical flow information. Metadata and configuration details, including sensor specifications and recording parameters, are saved in props.json. This structure facilitates straightforward access to each data type, improving data handling and compatibility with various processing pipelines. The directory structure is summarized below:
[…]
Figure 6: Structure of the data repository.
The official documentation provides detailed information on the contents of each HDF5 file. However, users can easily access or write data through the readers and writers included in eWiz without needing to handle these details directly.
7 Conclusion
This work addresses the challenge of developing reliable optical flow prediction networks for underwater environments, where collecting large-scale, annotated, event-based optical flow datasets is exceptionally difficult due to the complexities of underwater imaging. To overcome this, this work introduces the eStonefish-scenes dataset, along with a user-friendly pipeline to extend and adapt the dataset. To the best of our knowledge, no prior event-based datasets exist for underwater applications. The dataset incorporates scenes capturing event-based data from an ROV navigating a dynamic, coral-rich environment filled with diverse marine flora and fauna. We also introduce Stonefish-scenegen, an automated scene generator that eliminates the need for manually creating time-consuming reef environments, and Stonefish-boids, a package that simulates realistic schooling behaviour of fish. The dataset and supporting packages were created with the goal of accelerating research on EBCs for underwater applications, especially for training CNNs and SNNs. We also present eWiz, a comprehensive event-based data manipulation library that includes an optimized event data format. The library features data converters, writers, readers, and processing modules focused on data management, compression, and optimization. Additionally, it offers various loss functions, performance metrics, and visualizers to facilitate work with these datasets, along with an implementation of the contrast maximization algorithm. Both eWiz and eStonefish-scenes are designed to streamline research on EBCs and motion estimation for autonomous underwater vehicles by providing easy-to-use libraries for generating, processing, and evaluating event-based datasets.
{funding}
Jad Mansour was supported by the Joan Oró Grant no. 2024 FI-2 00762. The study was also supported in part by the Spanish government through the SIREC project no. PID2020-116736RB-IOO.
{dci}
The authors declared no potential conflicts of interest with respect to the research, authorship, and/or publication of this article.
{credit}
Jad Mansour: Conceptualization, Methodology, Software, Data Curation, Writing - original draft; Sebastian Realpe: Software, Writing - original draft; Hayat Rajani: Conceptualization, Methodology, Supervision, Validation, Writing - original draft; Michele Grimaldi: Software, Writing - review & editing; Rafael Garcia: Project Administration, Supervision, Writing - review & editing; Nuno Gracias: Project Administration, Supervision, Writing - review & editing.